OpenStudy (anonymous):

the distance between the two points -3 and 5

jigglypuff314 (jigglypuff314):

look at it this way 

jigglypuff314 (jigglypuff314):

distance means subtracting so 5 - (-3) = 5 + 3 = because subtracting a negative number means that you add the two numbers.

OpenStudy (anonymous):

Thanks